sedirg to a theory in genees ???ndcolorld plants ae aossed withshort andoololess plats, lour types of plants ? result td and calohdulind eolelna shot and calodd and iho inderknem w.cornpn tilies ol 16. 16. 16.and Click the con to view more indormation about a mulinomial experiment pot a mAhonilepein nt, ? 10 plants are selected td te prob bany that 3 will be al and colorM3?be tal and cokness ,w.besho tand oolert L ndrabehr-cams The probablity i 0 040 (Round to three decimal places as needed )
Answer #1

MultinonimalExperiment

P=\frac{n!}{n_1!n_2!n_3!.......n_k!}p_1^{n_1}p_2^{n_2} ......p_n^{n_k}

here in question

n1=3 ,p1=5/16
n2=3 ,p2=4/16
n3=1 , p3=1/16
n4=3 , p4=6/16

using all above values in formula

P=\frac{10!}{3!3!1!3!}(5/16)^{3}(4/16)^{3}(1/16)^1(6/16)^3

P=16800*0.03051*0.0156*0.05273*0.0625=0.02635

if you have any doubt ask in comment give thumbs up if you like work.